OpenStudy (anonymous):

mTo solve for y in the equation 2x + y = 5, subtract 2 from both sides of the equation.

jimthompson5910 (jim_thompson5910):

Close. It should be "subtract 2x from both sides"

jimthompson5910 (jim_thompson5910):

Doing that will give you 2x + y = 5 2x + y -2x= 5-2x y = -2x + 5

OpenStudy (anonymous):

true or false

jimthompson5910 (jim_thompson5910):

Well if it really says "2" and not "2x", then it's false. If it does say "2x", then it's true.

OpenStudy (anonymous):

The equation is equivalent to u = rst.

jimthompson5910 (jim_thompson5910):

What do you mean?

OpenStudy (anonymous):

true or false

jimthompson5910 (jim_thompson5910):

Something seems off

OpenStudy (anonymous):

The equation r= u/st is equivalent

OpenStudy (anonymous):

true or false

jimthompson5910 (jim_thompson5910):

which one do you think it is? and why?

OpenStudy (anonymous):

idk what it is

jimthompson5910 (jim_thompson5910):

how would you solve for u in r= u/st ?

OpenStudy (anonymous):

The expression 4y - 6 + 4y 2 - 9 contains three terms.

OpenStudy (anonymous):

The expression 4y - 6 + 4y 2 - 9 contains three terms.